OpenStudy (anonymous):

g(x) = 3x2 + 12x + 15 Put it in vertex form

OpenStudy (anonymous):

-12/ 2*15*3

OpenStudy (anonymous):

-2/15

OpenStudy (campbell_st):

well start by writing it as \[y = 3(x^2 + 4x) + 15\] complete the square on the terms inside the brackets.... what would you add for a perfect square?

OpenStudy (anonymous):

16???

OpenStudy (campbell_st):

not quite you add half of the coefficient of the term in x, and square it so (4/2)^2 = ?

OpenStudy (anonymous):

4

OpenStudy (campbell_st):

yes... so the equation is \[y = 3(x^2 + 4x + 4) + 15 - 3\times 4\] you just add 4 inside the brackets so to kee the equation in balance you need to subtract 3 lots of 4... so the equation is \[y = 3(x^2 + 4x + 4) + 3\] now just factor the quadratic, which is a perfect square to get the vertex form

OpenStudy (anonymous):

im still confused